question 1)what are the solutions to the equation x^2-6x+5=-8

Answer by mananth(16946)

x^2-6x+5=-8
..
x^2-6x+5+8=0
..
x^2-6x+13=0
..
discriminant = -16
..
x1= (( 6+sqrt(-16))/2
x1= 6/2 +sqrt(-16)/2
x1=3+(sqrt(16)*i)/2
x1=3+4i/2
x1=3+2i
...
x2= 3-2i
..
m.ananth@hotmail.ca